An extract from the Statement of Financial Position as at 31 March 2012 for Easy

FM Radio Station is as follows:

6% Secured Bonds (par $1000) $   400,000

Long Term Loans                       $   500,000

7% Preference Shares (par $2)                  $   250,000

Issued & Paid up ordinary shares ($1 par) $1,000,000

Retained Earnings $2,350,000

Other information:

The current market price for the bonds that will mature on 31 March 2019 is $875 each.

The interest rate on an equivalent long term loan available in the market today is 9%

The preference shares are currently trading at $1.95

The ordinary shares are currently trading at $2.65

The dividends per ordinary share have been as follows

2008 48 cents

2009 53 cents

2010 55 cents

2011 55 cents

2012 56 cents

It is expected that the dividend for 2013 will be 58 cents

The current tax rate for the firm is 28%.

Apply the cost of ordinary capital to retained earnings

Calculate the Weighted Average Cost of Capital (WACC) for Easy FM. Show your working
Answer #1

1.Calculation of Cost of 6% secured bonds:-

= Rate of interest (1- tax rate)

= 6(1-28%) = 4.32%

2. calculation of long term loans :-

= 9(1-28%) = 6.48%

3. cost of preference share-

= Preference dividend / market price per share *100

= .14 / 1.95*100

= 7.18%

4. cost of ordinary share :-

= Dividend per share / market price per share * 100

= .58 / 2.65*100

= 21.88%

Calculation of WACC:-

= Weight * Cost of respective source

= 4,00,000 / 45,00,000 * 4.32 +5,00,000/45,00,000*6.48 + 2,50,000/45,00,000*7.18 + 33,50,000/45,00,000*21.88

=0.38 + 7.2 + 0.40 + 16.29

= 24.27% WACC